OpenStudy (anonymous):

I don't know what to do with this question, please help! A piece of string 10m long is cut into pieces, so that the lengths of the pieces form an arithmetic sequence. (i) The lengths of the longest and shortes pieces are 1m and 25cm respectively; how many pieces are there? Answer:16 (ii) If the same string had been cut into 20 pieces with lengths that formed an arithmetic sequence, and if the length of the second longest had been 92.5 cm, how long would the shortest piece have been? (ii) Is the one I don't know what to do with. I have already completed (i)

OpenStudy (anonymous):

which one !?

OpenStudy (anonymous):

I need help with the last question please.

OpenStudy (welshfella):

Arithmetic Series: 2nd term = a1 + d where a = first term and d = common difference so here 92.5 = a1 + d

OpenStudy (welshfella):

the sum of 10 terms = the length of the string = 10,000 cms

OpenStudy (anonymous):

Um isn't 1000cm?

OpenStudy (welshfella):

Now the sum of n terms is (n/2)[2a1 + (n - 1)d] so here n = 20 and Sum = 10,000 so we have 10,000 = 10[2a1 + 19d]

OpenStudy (welshfella):

so we have the system of equations a1 + d = 92.5 20a1 + 19d = 10,000

OpenStudy (anonymous):

Ok, one little question- why is it 2a1 ?

OpenStudy (welshfella):

2 * first term a1 will be the first term = shortest piece

OpenStudy (anonymous):

Oh ok

OpenStudy (welshfella):

so you need to find the value of a1

OpenStudy (anonymous):

I use a different form of that equation normally so was confused for a second

OpenStudy (welshfella):

if you multiply the first equation by -19 and add you can find a1

OpenStudy (welshfella):

yes a1 is usually used in the US in UK we usually use just a

OpenStudy (anonymous):

a1 + d = 92.5 sorry, could you just explain more about how you got this part? I dont really understand why the first term plus the common term equals the second last term? I know I am missing something

OpenStudy (welshfella):

oh - maybe i was wrong there sorry - -i took that to be the second term - its the 19th term

OpenStudy (welshfella):

also the total length = 1000 (wrong again!)

OpenStudy (anonymous):

Haha, it happens! No worries.

OpenStudy (welshfella):

so the 2 equations are a1 + 18d = 92.5 20a1 + 19d = 1000

OpenStudy (anonymous):

Ok, I am not sure I understand the concept behind that second equation

OpenStudy (welshfella):

that is the formula for the sum of n terms of an arithmetic series Sn = (n/2)[2a1 + (n - 1)d] here Sn = length of the string = 1000 cms, n = 20 (20 pieces) , a1 = first term, d = common difference

OpenStudy (anonymous):

Ok. Then I am meant to simultaneously solve those two equations?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

i bit messy Have you got a graphical calculator?

OpenStudy (anonymous):

No I don't. Oh well, I will try it quickly now without one

OpenStudy (amistre64):

is our shortest piece still 25cm?

OpenStudy (anonymous):

No I don't think so. That was for the first question

OpenStudy (amistre64):

how long should shortest piece ... forgot to read thru it all the way

OpenStudy (welshfella):

hmm the results don't seem to give rational values..

OpenStudy (anonymous):

Yeah I am not getting anything that makes sense.

OpenStudy (anonymous):

I got d=2.5 a=47.625 after rounding

OpenStudy (welshfella):

i thnk the equations are correct..

OpenStudy (anonymous):

I probably just did something wrong then.

OpenStudy (welshfella):

yes the calc gives 47.631 and 2.4927

OpenStudy (amistre64):

10m long 20(2a+19d)/2 = 1000 a+18d = 92.5 a is about 2.5, and d is about 5

OpenStudy (welshfella):

yea my second equation was wrong 10(2a + 19d) = 201 + 190d

OpenStudy (amistre64):

2a+19d = 100 a+18d = 92.5 2a +19d = 100 -19/18a -19d = -19/18(92.5) a = (100-(19/18(92+1/2)))/(2-(19/18))

OpenStudy (welshfella):

2.5 for shortest length and 5 for common difference

OpenStudy (welshfella):

it was 190d not 19d

OpenStudy (anonymous):

Ok..

OpenStudy (welshfella):

do you see the mistake I didnt distribute 10 over the (2a + 19) correctly

OpenStudy (welshfella):

I thought someone would spot my deliberate mistake - oh amistre did!

OpenStudy (anonymous):

Oh! I see now!

OpenStudy (welshfella):

good there is another formula for sum of n terms of an arithmetic series Sn = (n/2)[ a1 + L] where L is the last term If you know the last term its a lot simpler

OpenStudy (anonymous):

Yes, thats the form I normally use and I didn't remeber about the other one not needing the last term which is a big part of why I got stuck I think Just trying to solve those two equations myself now.

OpenStudy (welshfella):

probably best to multiply first one by -20 and add to get d first

OpenStudy (anonymous):

Yup! Did it and got the two right answers! Thank you so much for your help!

OpenStudy (welshfella):

yw